# On combinatorial invariance of parabolic Kazhdan–Lusztig polynomials

### v1

@@ 1. Introduction
Let $[u,v]^{J}\coloneqq[u,v]\cap W^{J}$ denote the set of elements from $[u,v]$ lying in $W^{J}$. Brenti, in the case of Hermitian symmetric pairs [6, Corollary 4.8], proved that parabolic Kazhdan–Lusztig polynomials are equal when $[u_{1},v_{1}]^{J}\cong[u_{2},v_{2}]^{J}$ (see also [4, 13]). But examples from [8, 14] show that the naïve extensions of Conjecture 1.1 that this suggests, where one only requires that $[u_{1},v_{1}]^{J_{1}}\cong[u_{2},v_{2}]^{J_{2}}$ or that $[u_{1},v_{1}]\cong[u_{2},v_{2}]$, are false in general. Thus any extension to the parabolic setting must include information about the isomorphism type of the Bruhat intervals as well as information about their intersections with $W^{J}$. Marietti proposed such an extension and proved it for lower intervals.
@@
Conjecture 1.3(a) implies Conjecture 1.2 since any poset isomorphism $[u_{1},v_{1}]^{J_{1}}\to[u_{2},v_{2}]^{J_{2}}$ in particular restricts to a bijection $A^{J_{1}}_{u_{1},v_{1}}\to A^{J_{2}}_{u_{2},v_{2}}$. Conjecture 1.2 in turn implies Conjecture 1.1 by taking $J_{1}=J_{2}=\emptyset$. In our first main result, we show that the three conjectures are in fact equivalent. The equivalence of Conjectures 1.1 and 1.2 is already new.
@@
Conjectures 1.1, 1.2 and 1.3 hold in the case $u_{1}=e_{1}$ and $u_{2}=e_{2}$ are the identity elements of $W_{1}$ and $W_{2}$.
@@
Conjectures 1.1, 1.2 and 1.3 hold when $[u_{1},v_{1}]$ is a short edge interval.
@@
Section 2 gives background on Bruhat order and Kazhdan–Lusztig polynomials. In Section 3 we introduce invariant collections of Bruhat intervals and prove Theorem 3.3 which gives general criteria the for combinatorial invariance of parabolic Kazhdan–Lusztig polynomials. Theorems 1.5, 1.6, 1.7 and 1.8 will be shown in Section 4 to follow from Theorem 3.3 and known special cases of Conjecture 1.1.
@@ 3. Invariant collections of intervals
Theorem 3.3 below is our most general result. Theorems 1.5, 1.6, 1.7 and 1.8 will all be shown to follow from it.
@@
If moreover $\varphi$ sends $[u_{1},v_{1}]^{J_{1}}$ to $[u_{2},v_{2}]^{J_{2}}$, then:
@@
    Suppose now that $\varphi$ sends $[u_{1},v_{1}]^{J_{1}}$ to $[u_{2},v_{2}]^{J_{2}}$ and that $\mathcal{I}$ is fully invariant. For any $u_{1}^{\prime},v_{1}^{\prime}\in[u_{1},v_{1}]^{J_{1}}$ the restriction $\varphi|_{[u_{1}^{\prime},v_{1}^{\prime}]}$ is a poset isomorphism onto its image $[u_{2}^{\prime},v_{2}^{\prime}]$, where $u_{2}^{\prime}\coloneqq\varphi(u_{1}^{\prime}),v_{2}^{\prime}\coloneqq\varphi(v_{1}^{\prime})$. Furthermore, it sends $A^{J_{1}}_{u_{1}^{\prime},v_{1}^{\prime}}$ to $A^{J_{2}}_{u_{2}^{\prime},v_{2}^{\prime}}$. Thus, by the arguments in part (a), we have that
@@
    (6) $q^{\ell(v_{1})-\ell(u_{1})}P^{J_{1},x}_{u_{1},v_{1}}(q^{-1})-P^{J_{1},x}_{u_{1},v_{1}}(q)$ $=\sum_{\begin{subarray}{c}\sigma_{1}\in W_{1}^{J_{1}}\\ u_{1}<\sigma_{1}\leq v_{1}\end{subarray}}R^{J_{1},x}_{u_{1},\sigma_{1}}(q)P^{J_{1},x}_{\sigma_{1},v_{1}}(q)$
    (9) $=\sum_{\begin{subarray}{c}\sigma_{2}\in W_{2}^{J_{2}}\\ u_{2}<\sigma_{2}\leq v_{2}\end{subarray}}R^{J_{2},x}_{u_{2},\sigma_{2}}(q)P^{J_{2},x}_{\sigma_{2},v_{2}}(q)$
@@
It follows from [7, Thm. 7.8] that the collection of lower intervals in $W_{1}$ and $W_{2}$ is fully invariant. Applying Theorem 3.3(a) and (b) proves Conjecture 1.3 (a) and (b) for lower intervals, and applying Theorem 3.3(c) yields this case of Conjecture 1.2. ∎
@@
It follows from [5, Thm. 6.3] that the collection $\mathcal{I}$ of short edge intervals in $W_{1}$ and $W_{2}$ is fully invariant. Applying Theorem 3.3(a) and (b) proves Conjecture 1.3 (a) and (b) for intervals from $\mathcal{I}$, and applying Theorem 3.3(c) yields this case of Conjecture 1.2. ∎
@@ Acknowledgements
We are grateful to Mario Marietti for his helpful comments.

### v2

@@ 1. Introduction
Conjecture 1.1 has received considerable study, with many special cases having been established. In particular, it has been proven for lower intervals [8] and short edge intervals [6]. It has also been verified for Coxeter groups of type $\widetilde{A}_{2}$ [11] and for small rank finite Coxeter groups [14]. Work on this conjecture in the case of type $A$ has been especially active in recent years thanks to the conjectural approach of [4].
Remark 1.2.
@@
Let $[u,v]^{J}\coloneqq[u,v]\cap W^{J}$ denote the set of elements from $[u,v]$ lying in $W^{J}$. Brenti, in the case of Hermitian symmetric pairs [7, Corollary 4.8], proved that parabolic Kazhdan–Lusztig polynomials are equal when $[u_{1},v_{1}]^{J}\cong[u_{2},v_{2}]^{J}$ (see also [5, 16]). But examples from [9, 17] show that the naïve extensions of Conjecture 1.1 that this suggests, where one only requires that $[u_{1},v_{1}]^{J_{1}}\cong[u_{2},v_{2}]^{J_{2}}$ or that $[u_{1},v_{1}]\cong[u_{2},v_{2}]$, are false in general. Thus any extension to the parabolic setting must include information about the isomorphism type of the Bruhat intervals as well as information about their intersections with $W^{J}$. Marietti proposed such an extension and proved it for lower intervals.
@@
Conjecture 1.4(a) implies Conjecture 1.3 since any poset isomorphism $[u_{1},v_{1}]^{J_{1}}\to[u_{2},v_{2}]^{J_{2}}$ in particular restricts to a bijection $A^{J_{1}}_{u_{1},v_{1}}\to A^{J_{2}}_{u_{2},v_{2}}$. Conjecture 1.3 in turn implies Conjecture 1.1 by taking $J_{1}=J_{2}=\emptyset$. In our first main result, we show that the three conjectures are in fact equivalent. The equivalence of Conjectures 1.1 and 1.3 is already new.
@@
Conjectures 1.1, 1.3 and 1.4 hold in the case $u_{1}=e_{1}$ and $u_{2}=e_{2}$ are the identity elements of $W_{1}$ and $W_{2}$.
@@
Conjectures 1.1, 1.3 and 1.4 hold when $[u_{1},v_{1}]$ is a short edge interval.
@@
Section 2 gives background on Bruhat order and Kazhdan–Lusztig polynomials. In Section 3 we introduce invariant collections of Bruhat intervals and prove Theorem 3.3 which gives general criteria the for combinatorial invariance of parabolic Kazhdan–Lusztig polynomials. Theorems 1.6, 1.7, 1.8 and 1.9 will be shown in Section 4 to follow from Theorem 3.3 and known special cases of Conjecture 1.1.
@@ 3. Invariant collections of intervals
Theorem 3.3 below is our most general result. Theorems 1.6, 1.7, 1.8 and 1.9 will all be shown to follow from it.
@@
If moreover $\varphi$ restricts to an isomorphism $[u_{1},v_{1}]^{J_{1}}\to[u_{2},v_{2}]^{J_{2}}$, then:
@@
    Suppose now that $\varphi$ restricts to an isomorphism $[u_{1},v_{1}]^{J_{1}}\to[u_{2},v_{2}]^{J_{2}}$ and that $\mathcal{I}$ is fully invariant. For any $u_{1}^{\prime},v_{1}^{\prime}\in[u_{1},v_{1}]^{J_{1}}$ the restriction $\varphi|_{[u_{1}^{\prime},v_{1}^{\prime}]}$ is a poset isomorphism onto its image $[u_{2}^{\prime},v_{2}^{\prime}]$, where $u_{2}^{\prime}\coloneqq\varphi(u_{1}^{\prime}),v_{2}^{\prime}\coloneqq\varphi(v_{1}^{\prime})$. Furthermore, it sends $A^{J_{1}}_{u_{1}^{\prime},v_{1}^{\prime}}$ to $A^{J_{2}}_{u_{2}^{\prime},v_{2}^{\prime}}$. Thus, by the arguments in part (a), we have that
@@
    $q^{\ell(v_{1})-\ell(u_{1})}P^{J_{1},x}_{u_{1},v_{1}}(q^{-1})-P^{J_{1},x}_{u_{1},v_{1}}(q)$ $=\sum_{\begin{subarray}{c}\sigma_{1}\in W_{1}^{J_{1}}\\ u_{1}<\sigma_{1}\leq v_{1}\end{subarray}}R^{J_{1},x}_{u_{1},\sigma_{1}}(q)P^{J_{1},x}_{\sigma_{1},v_{1}}(q)$
    $=\sum_{\begin{subarray}{c}\sigma_{2}\in W_{2}^{J_{2}}\\ u_{2}<\sigma_{2}\leq v_{2}\end{subarray}}R^{J_{2},x}_{u_{2},\sigma_{2}}(q)P^{J_{2},x}_{\sigma_{2},v_{2}}(q)$
@@
It follows from [8, Thm. 7.8] that the collection of lower intervals in $W_{1}$ and $W_{2}$ is fully invariant. Applying Theorem 3.3(a) and (b) proves Conjecture 1.4 (a) and (b) for lower intervals, and applying Theorem 3.3(c) yields this case of Conjecture 1.3. ∎
@@
It follows from [6, Thm. 6.3] that the collection $\mathcal{I}$ of short edge intervals in $W_{1}$ and $W_{2}$ is fully invariant. Applying Theorem 3.3(a) and (b) proves Conjecture 1.4 (a) and (b) for intervals from $\mathcal{I}$, and applying Theorem 3.3(c) yields this case of Conjecture 1.3. ∎
@@ Acknowledgements
We are grateful to Mario Marietti and Paolo Sentinelli for interesting discussions during the Workshop “Bruhat order: recent developments and open problems” held at the University of Bologna.